A metal bar 50cm long at 15°c is heated to 85°c. If it expands by 0.088cm, determine its linear expansivity.

1
Expert's answer
2022-02-18T08:21:33-0500

Given:

"l_0=50\:\rm cm"

"\Delta l=0.088\:\rm cm"

"\Delta t=70\:\rm K"


The thermal expansion is described by an equation

"\Delta l=l_0\alpha\Delta t"

Hence, the linear expansivity

"\alpha=\Delta l/(l_0\Delta t)"

"\alpha=\rm 0.088\: cm/(50\: cm*70 \: K)=2.1*10^{-5}\: K^{-1}"